OpenStudy (anonymous):

integrate dx / ((x^2) (sqrt(25-x^2))) using trig substitutions?

OpenStudy (anonymous):

i think you use \(x=5\sin(u)\) so that \[\sqrt{25-x^2}=\cos(u)\]

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

\[\int\sin^2(u)^2\cos^2(u)du\]

OpenStudy (anonymous):

what is u?

OpenStudy (anonymous):

if you make the substitution \[x=5\sin(u)\] the \[u=\sin^{-1}(\frac{x}{5})\]

OpenStudy (anonymous):

okay so what i did so far is replace x = 5sin u and dx=5cosu du int 5cosu du/ 5sinu sqrt(25-(5sin^2 u)) sub y = sin u and dy = cosu du dy/y^2 (5-u)

OpenStudy (anonymous):